# The Impact of LLMs on Online News Consumption and Production

### v2

@@ 4.4 Effect 4: Publishers Shift Toward Media-Rich Content Rather Than Scaling Text
We next examine how publishers adjust their content-supply strategies. Using the HTTP Archive, we measure webpage composition by counting specific HTML elements across functional categories: article volume (e.g., <article>, <section>), visual and multimedia richness (e.g., <img>, <video>, <figure>), interactive mechanisms (e.g., <button>, <input>), and advertising and targeting technologies (e.g., <script>, <iframe>). To control for broader web development trends, we compare these patterns with a control group of the top 100 retail domains. Figure 7(f) plots time series of average element counts for each category for both the publisher and retail sectors, while Figures A6 through A9 present the corresponding disaggregated element counts.
@@
Using the Internet Archive’s Wayback Machine, we collect annual counts of newly observed URLs by content type. Figure A10 plots the number of unique URLs of different types, which is consistent with findings from the HTTP Archive’s element counts. We do not find the number of new text/article URLs to be increasing over the years; rather, there is a moderate decrease, suggesting that publishers are not responding by producing more articles. Instead, growth is concentrated in non-text assets: the number of unique image URLs increases substantially. Together, these patterns imply that publishers’ primary adjustment margin is to enrich existing content with additional media and embedded components rather than scaling up text output.
@@ 5 Discussion and Conclusion
We study how generative AI is reshaping the newspaper publishing industry by linking high-frequency measures of audience demand, access control, organizational hiring, and content format. Across multiple data sources, we document four effects. First, total visits to publishers are broadly stable until mid-2024, with a noticeable and persistent decline emerging only after August 2024. Second, publishers adopt GenAI crawler blocking in a staggered pattern starting in mid-2023, and staggered DiD estimates suggest that blocking is followed by a persistent decline in total traffic; importantly, the decline also appears in Comscore-based human browsing, indicating economically meaningful audience effects rather than a purely mechanical removal of bot activity for large publishers. Third, publisher hiring does not exhibit a disproportionate contraction in core editorial/content roles relative to other roles, suggesting limited near-term labor substitution away from newsroom production. Fourth, publishers do not scale text output; instead, they shift toward richer pages with more interactive components and greater use of advertising and targeting technologies, with growth concentrated in image-related URLs and interactive elements. Together, these facts suggest that generative AI is not yet functioning as a complete substitute for traditional news production and monetization; rather, publishers appear to be adjusting along multiple domains: access control, hiring composition, and content format and quantity production.
@@
Several limitations qualify interpretation. Our traffic measures combine modeled aggregates (SimilarWeb) with a U.S. desktop panel (Comscore), and neither directly observes LLM-mediated consumption paths (e.g., in-chat answers, referrals, or citation links). Robots.txt measures stated restrictions rather than enforceable access control, and blocking decisions may coincide with other time-varying publisher actions (paywall changes, SEO strategy, site redesigns, or platform shocks). Finally, the study covers an early phase of the technology, before newer interfaces and integration (e.g., more prominent AI summaries and search-chat products) may fully diffuse.
@@
Table A1: Distribution of URLs by NAICS Sector
@@
Table A2: Website Traffic in 2024 SimilarWeb Data
@@
Table A3: AR(1) Break Regressions at Detected Cutoff Dates
@@
Table A4: AR(1) Break Regressions at Placebo Cutoff Dates
@@
Table A5: Major GenAI / LLM-related crawlers and user agents
@@
Figure A1: News Publishing Website Traffic (August 2024)
@@
This table reports the average treatment effect on the treated (ATT) for website traffic following August 2024. The dependent variable is the logarithm of daily website traffic ($\log(\text{Traffic})$). The control group consists of the top 100 retail websites ranked by daily traffic. The left panel plots the Synthetic Difference-in-Differences (SDID) estimator, while the right panel shows the standard TWFE DiD estimator. Standard errors are calculated using 50 bootstrap repetitions for SDID and robust standard errors clustered at the publisher level for TWFE DiD.
Figure A2: News Publishing Website Traffic (November 2023)
@@
This table reports the average treatment effect on the treated (ATT) for website traffic following November 2023. The dependent variable is the logarithm of daily website traffic ($\log(\text{Traffic})$). The control group consists of the top 100 retail websites ranked by daily traffic. The left panel plots the Synthetic Difference-in-Differences (SDID) estimator, while the right panel shows the standard TWFE DiD estimator. Standard errors are calculated using 50 bootstrap repetitions for SDID and robust standard errors clustered at the publisher level for TWFE DiD.
Figure A3: News Publishing Website Traffic (April 2023)
@@
This table reports the average treatment effect on the treated (ATT) for website traffic following the April 2023 treatment. The dependent variable is the logarithm of daily website traffic ($\log(\text{Traffic})$). The control group consists of the top 100 retail websites ranked by daily traffic. The left panel plots the Synthetic Difference-in-Differences (SDID) estimator, while the right panel shows the standard TWFE DiD estimator. Standard errors are calculated using 50 bootstrap repetitions for SDID and robust standard errors clustered at the publisher level for TWFE DiD.
Table A6: Synthetic and TWFE DiD estimates of blocking GenAI crawlers on publisher traffic
Log(SimilarWeb traffic) Log(Comscore traffic)
SDID TWFE DiD SDID TWFE DiD
ATT -0.179*** -0.179*** -0.199*** -0.199***
[-0.273,-0.085] [-0.279,-0.079] [-0.258,-0.139] [-0.252,-0.145]
This table reports Synthetic and TWFE DiD ATT estimates of the effect of blocking GenAI web crawlers on publisher traffic. The dependent variables in columns (1)–(2) and (3)–(4) are the logarithm of monthly visits from SimilarWeb and Comscore, respectively. Confidence intervals are based on 50 bootstrap repetitions. *p<<0.1; **p<<0.05; ***p<<0.01.
Figure A4: GenAI Bot Blocking Fractions by News Publishers Traffic Rank Group and Retailers
@@
Figure A5: Site Framework and Layout
Figure A6: Article Volume
Figure A7: Interactive Features
Figure A8: Visual & Media Richness
Figure A9: Ads & Targeting
Figure A10: Internet Archive’s Wayback Machine Content Type Unique URLs

### v3

@@ 4.4 Effect 4: Publishers Shift Toward Media-Rich Content Rather Than Scaling Text
We next examine how publishers adjust their content-supply strategies. Using the HTTP Archive, we measure webpage composition by counting specific HTML elements across functional categories: article volume (e.g., <article>, <section>), visual and multimedia richness (e.g., <img>, <video>, <figure>), interactive mechanisms (e.g., <button>, <input>), and advertising and targeting technologies (e.g., <script>, <iframe>). To control for broader web development trends, we compare these patterns with a control group of the top 100 retail domains. Figure 8(f) plots time series of average element counts for each category for both the publisher and retail sectors, while Figures 17 through 20 present the corresponding disaggregated element counts.
@@
Using the Internet Archive’s Wayback Machine, we collect annual counts of newly observed URLs by content type. Figure 21 plots the number of unique URLs of different types, which is consistent with findings from the HTTP Archive’s element counts. We do not find the number of new text/article URLs to be increasing over the years; rather, there is a moderate decrease, suggesting that publishers are not responding by producing more articles. Instead, growth is concentrated in non-text assets: the number of unique image URLs increases substantially. Together, these patterns imply that publishers’ primary adjustment margin is to enrich existing content with additional media and embedded components rather than scaling up text output.
@@ 5 Discussion and Conclusion
We study how generative AI is reshaping the newspaper publishing industry by linking high-frequency measures of audience demand, access control, organizational hiring, and content format. Across multiple data sources, we document four effects. First, total visits to publishers are broadly stable until mid-2024, with a noticeable and persistent decline emerging only after August 2024. Second, publishers blocked GenAI crawler in a staggered pattern starting in mid-2023, and staggered DiD estimates suggest a significant decline in traffic in the first few months following adoption. We also observe a noisy decline in Comscore-based human browsing, suggesting potentially meaningful audience effects rather than a purely mechanical removal of bot activity for large publishers. Third, publisher hiring does not exhibit a disproportionate contraction in core editorial/content roles relative to other roles, suggesting limited near-term labor substitution away from newsroom production. Fourth, publishers do not scale text output; instead, they shift toward richer pages with more interactive components and greater use of advertising and targeting technologies, with growth concentrated in image-related URLs and interactive elements. Together, these facts suggest that generative AI is not yet functioning as a complete substitute for traditional news production and monetization; rather, publishers appear to be adjusting along multiple domains: access control, hiring composition, and content format and quantity production.
@@
Several limitations qualify interpretation. Our traffic measures combine modeled aggregates (SimilarWeb) with a U.S. desktop panel (Comscore), and neither directly observes LLM-mediated consumption paths (e.g., in-chat answers, referrals, or citation links). Robots.txt measures stated restrictions rather than enforceable access control. Finally, the study covers an early phase of the technology, before newer interfaces and integration (e.g., more prominent AI summaries and search-chat products) may fully diffuse.
@@
Table 8: Distribution of URLs by NAICS Sector
@@
Table 9: Website Traffic in 2024 SimilarWeb Data
@@
Table 10: AR(1) Break Regressions at Detected Cutoff Dates
@@
Table 11: AR(1) Break Regressions at Placebo Cutoff Dates
@@
Figure 9: Total, Direct, and Google Referral weekly traffic trends.
The plot compares Total Traffic against specific referral sources (Google.com and Direct) from Comscore data. The vertical dashed line indicates July 1, 2024.
Figure 10: Weekly traffic comparison for Direct vs Search Referral traffic
(a) US Mobile Traffic
(b) US Desktop Traffic
The plots compare Direct traffic versus Organic Search referrals on Mobile (a) and Desktop (b) from Semrush data. The vertical dashed line indicates July 1, 2024.
Table 12: Major GenAI / LLM-related crawlers and user agents
@@
Figure 11: News Publishing Website Traffic (August 2024)
@@
This figure reports the average treatment effect on the treated (ATT) for website traffic following August 2024. The dependent variable is the logarithm of daily website traffic ($\log(\text{Traffic})$). The control group consists of the top 100 retail websites ranked by daily traffic. The left panel plots the Synthetic Difference-in-Differences (SDID) estimator, while the right panel shows the standard TWFE DiD estimator. Standard errors are calculated using 50 bootstrap repetitions for SDID and robust standard errors clustered at the publisher level for TWFE DiD.
Figure 12: News Publishing Website Traffic (November 2023)
@@
This figure reports the average treatment effect on the treated (ATT) for website traffic following November 2023. The dependent variable is the logarithm of daily website traffic ($\log(\text{Traffic})$). The control group consists of the top 100 retail websites ranked by daily traffic. The left panel plots the Synthetic Difference-in-Differences (SDID) estimator, while the right panel shows the standard TWFE DiD estimator. Standard errors are calculated using 50 bootstrap repetitions for SDID and robust standard errors clustered at the publisher level for TWFE DiD.
Figure 13: News Publishing Website Traffic (April 2023)
@@
This figure reports the average treatment effect on the treated (ATT) for website traffic following the April 2023 treatment. The dependent variable is the logarithm of daily website traffic ($\log(\text{Traffic})$). The control group consists of the top 100 retail websites ranked by daily traffic. The left panel plots the Synthetic Difference-in-Differences (SDID) estimator, while the right panel shows the standard TWFE DiD estimator. Standard errors are calculated using 50 bootstrap repetitions for SDID and robust standard errors clustered at the publisher level for TWFE DiD.
Figure 14: GenAI Bot Blocking Fractions by News Publishers Traffic Rank Group and Retailers
@@
This figure plots the fraction of websites that disallow GenAI bots over time. Panel (a), Group 1, includes the largest publishers (more than 10 visits per day on average); Panel (b), Group 2, includes mid-sized publishers (1–10 visits per day; ranked 34th–164th by average daily visits); Panel (c), Group 3, includes the smallest publishers (fewer than 1 visit per day on average); and Panel (d) includes the top 100 retailers.
Figure 15: Staggered DiD of blocking GenAI bots on publisher traffic (extended window)
This figure reports staggered DiD event-study estimates of the effect of blocking GenAI web crawlers on publisher traffic, using a 24-week window before and after blocking. The outcome in each panel is the logarithm of weekly visits from SimilarWeb traffic. Confidence intervals are based on 50 bootstrap replications.
Table 13: Lagged traffic and blocking adoption
Block
$\Delta\log(\text{Traffic})_{t-1}$ 0.00591
Constant -0.0790
(-1.02)
Week fixed effects Yes
The dependent variable is an indicator for first-time blocking adoption in week $t$ (at-risk sample). The regressor is lagged weekly traffic growth, $\Delta\log(\text{Traffic})_{t-1}$. Week fixed effects are included. t-statistics are in parentheses. *p<<0.1; **p<<0.05; ***p<<0.01.
Figure 16: Site Framework and Layout
Figure 17: Article Volume
Figure 18: Interactive Features
Figure 19: Visual & Media Richness
Figure 20: Ads & Targeting
Figure 21: Internet Archive’s Wayback Machine Content Type Unique URLs